You have an upstream Picking department that feeds two downstream Packing departments: Pack Singles and Pack Multis. Those Packing departments feed to a Shipping department that loads the outgoing trucks. DEPARTMENT % OF VOLUME RATE (UPH) VOLUME HEADCOUNT SINGLES 40% 104 MULTIS 60% 215 40% of your Pick volume goes to Pack Singles and has a packing rate of 104 units per labor hour (uph). 60% of the Pick volume goes to Pack Multis and has a pack rate of 215 units per labor hour. - Your pickers pick both Single and Multi items throughout the day at an overall average rate of 114 units per labor hour. All units that are packed in both processes go through the Ship process at a rate of 570 units per hour. - You have 102 people today for all 4 departments and you absolutely must pack 47,880 units in Pack Multis Items to meet a customer promise metric. - The Ops manager for the shift started building the plan but was called away. You have to finish it. He correctly allocated labor to the Pack Singles before he was called away. How do you allocate the remaining labor to balance the flow in your department if you work 10 hour shift? Do not assume breaks or lunches in your answer.
Answer #1

We have to pack 47880 units in Multis which is accounted for 60% of the volume.

Hence, Overall volume = 47880/60% = 79800

Volume of Single = 79800 - 47880 = 31920

For Picker -

Total Volume - 79800

Picking Rate = 114

Required Number of Labor Hours = 79800/114 = 700

Total Labor Required = 700/10 = 70 labor

For Multis -

Total Volume - 47880

Rate = 215

Required Number of Labor Hours = 47880/215 = 222.70

Total Labor Required = 222.70/10 = 22.27 = 23 labor

For Single -

Total Volume - 31920

Rate = 104

Required Number of Labor Hours = 31920/104 = 306.92

Total Labor Required = 306.92/10 = 30.69 = 31 labor

For Shipping -

Total Volume - 79800

Rate = 570

Required Number of Labor Hours = 79800/570 = 140

Total Labor Required = 140/10 = 14 labor

Total Labor Required as per given criteria = 70+23+31+14 = 138.

Hence, it will not be possible to attain the required flow with 102 labor with given criteria.